OpenStudy (anonymous):

A is now twice as old as B was when A was as old as B is now. When B is as old as A is now, their combine age s will be 63. How old are A and B now?

OpenStudy (anonymous):

Let the present ages of A and B be x and y respectively then, The difference in age between A and B is x-y...... Now, x=2(y-(x-y)) x=4y-2x 3x=4y.........(i)

OpenStudy (anonymous):

y+(x-y)+x+(x-y)=63 3x-y=63 y=3x-63....(ii)

OpenStudy (anonymous):

Solve equation (i) and (ii)...

OpenStudy (anonymous):

Got it?

OpenStudy (anonymous):

Yeah thanks

OpenStudy (anonymous):

welcome